### v4

# Strong and weak wave turbulence regimes in Bose-Einstein condensates

Abstract
When a turbulent Bose-Einstein condensate is driven out-of-equilibrium at a scale much smaller than the system size, nonlinear wave interactions transfer particles towards large scales in an inverse cascade process. In this work, we study numerically wave turbulence in a three-dimensional Bose-Einstein condensate in forced and dissipated inverse cascade settings. We observe that when the forcing rate increases, thereby increasing the particle flux, the turbulence spectrum gradually transitions from the weak-wave Kolmogorov-Zakharov cascade to a critical balance state characterized by a range of scales with balanced linear and nonlinear dynamic timescales. Further forcing increases lead to a coherent condensate component superimposed with Bogoliubov-type acoustic turbulence. The role of vortices in such a strongly forced state is marginal, which makes this new state very different from the strongly turbulent state composed from a tangle of quantized vortex lines. We then use our predictions and numerical data to formulate a new out-of-equilibrium equation of state for the $3D$ inverse cascade.
preprint: AIP/123-QED
Introduction.— Turbulence is a central example of strongly non-equilibrium systems whose state is determined by an energy (or another invariant) flux through scales. When forcing and dissipation scales are well separated, an inertial range of scales emerges, where the dynamics is driven by the nonlinear interactions only. The central concept of turbulence is the idea of cascade, originally stated by Richardson in the case of hydrodynamic turbulence. In 1941, Kolmogorov pushed this concept further by postulating that in the limit of an infinite inertial range, the energy dissipation rate $\varepsilon$ is finite, necessarily equal to the energy flux, and it is the only dimensional parameter involved in the turbulent energy cascade. A simple dimensional analysis is then enough to fix scaling exponents, which leads to the famous Kolmogorov energy spectrum $E_{k}\propto\varepsilon^{2/3}k^{-5/3}$, with $k$ the wavevector Frisch (1995). Many other turbulent systems exist where more than one dimensional parameter comes into play, so scaling exponents cannot be fixed so simply. Most typical examples are magnetohydrodynamics and nonlinear wave systems such as internal and inertial waves in the ocean and atmospheres, capillary-gravity waves in the surface of liquids, or Bose-Einstein condensates GALTIER et al. (2000); Galtier (2003); Lvov and Tabak (2001); Falcon and Mordant (2022). Fortunately, when nonlinearity is small, the weak wave turbulence theory (WWT) predicts the whole spectrum, including universal dimensionless pre-factors Nazarenko (2011). When nonlinearity becomes strong, only phenomenological arguments can be given in some scarce cases. How a system transits from the weak to the strong wave turbulence regimes remains an open question.
In this Letter we study the weak to strong transition in turbulence of Bose-Einstein condensates (BECs). BECs provide a great experimental platform for studying turbulence in general and WT in particular because of the versatility of optical techniques in creating experimental setups and diagnostics. Recent experimental groups have succeeded in creating turbulent BECs in two and three dimensions, by shaking the BEC trap and providing a synthetic dissipation mechanism Navon et al. (2016, 2019); Gałka et al. (2022); Karailiev et al. (2024). To model BECs, we use the Gross-Pitaevski equation (GPE) for the complex wave function $\psi({\bf x},t)$,
$i\frac{\partial\psi}{\partial t}=-\alpha\nabla^{2}\psi+\beta|\psi|^{2}\psi,$ (1)
where $\alpha={\hbar}/{2m}$ and $\beta=g/\hbar$, with $\hbar$ the reduced Plank constant, $m$ the boson mass and $g$ the interaction constant. For simplicity, we consider a triply periodic cube of side $L$ and volume $V=L^{3}$. Then the GPE (1) conserves the total number of particles and energy per unit volume,
$N$ $=$ $\frac{1}{L^{3}}\int|\psi(\mathbf{x},t)|^{2}\,\mathrm{d}\,\mathbf{x},$ (2)
$H$ $=$ $\frac{2\alpha}{L^{3}}\int\left[\alpha|\nabla\psi(\mathbf{x},t)|^{2}+\frac{\beta}{2}|\psi(\mathbf{x},t)|^{4}\right]\,\mathrm{d}\,\mathbf{x}\,$, (3)
respectively. In a turbulent setting, energy is transferred towards small scales (direct cascade), whereas particles exhibit an inverse cascade towards large scales. Most of turbulent BEC experiments have studied the direct energy cascade Navon et al. (2019); Gałka et al. (2022), whereas inverse particle turbulent transfers have been recently observed both in 2D and 3D experiments Karailiev et al. (2024); Moreno-Armijos et al. (2025). (The inverse particle cascade should not be confused with the inverse energy cascades and clustering in the 2D BEC quantized vortex turbulence studied in Johnstone et al. (2019); Gauthier et al. (2019).)
In this Letter, we focus on the inverse particle cascade in 3D. In particular, we study how the system transitions from weak to strong wave turbulence when the particle flux per unit of volume $|Q_{0}|$ increases. The advantage of this setting is that the WWT provides an exact and simple analytical prediction, unlike the 3D direct energy cascade or the cascades in 2D BECs that present some mathematical issues.
As usual in turbulence, let us start by dimensional considerations. The main observable in the turbulent cascade is the waveaction spectrum, $n_{\bf k}(t)\equiv n({\bf k},t)=\frac{V}{(2\pi)^{3}}\langle|\hat{\psi}_{\bf k}(t)|^{2}\rangle,$ where $\hat{\psi}_{\bf k}=\frac{1}{L^{3}}\int\psi(\mathbf{x})e^{i{\bf{k}}\cdot\mathbf{x},}\,\mathrm{d}\,\mathbf{x}$ is the Fourier transform of $\psi$, and the brackets denote the average over fluctuations. Note that $n_{\bf k}$ is dimensionless, and from Eq. (1), $\alpha]=length^{2}/time$ and $\beta]=length^{3}/time$. As $Q_{0}]=length^{-3}/time$, then by matching dimensions we have
$n_{\bf k}=C\left(\frac{|Q_{0}|\beta^{5}}{\alpha^{6}}\right)^{y}\left(\frac{\alpha}{\beta k}\right)^{x},$ (4)
where $C$ is a dimensionless constant, the two terms in parentheses are independent dimensionless combinations and we assumed a power law dependence on these parameters with $x$ and $y$ being a priori arbitrary exponents. Here, $x$ fixes the dependence of the waveaction spectrum on scales whereas $y$ determines the behavior of the so-called out-of-equilibrium equation of state Dogra et al. (2023); Zhu et al. (2024). Indeed, the the spectrum’s amplitude dependence on the flux $Q_{0}$ is analogous to the particle number dependence on the chemical potential at thermal equilibrium. The WWT predicts the Kolmogorov-Zakharov (KZ) spectrum that fixes all the exponents, and the universal dimensionless constant $C$ Zhu et al. (2023):
$n_{k}=C_{\rm i}\frac{\alpha^{1/3}}{\beta^{2/3}}|Q_{0}|^{1/3}k^{-7/3}\,,\quad C_{\rm i}\approx 7.5774045\times 10^{-2}.$ (5)
This spectrum is an exact solution of the GP-associated wave kinetic equation, and it was found to be in excellent agreement with numerical simulations Zhu et al. (2023).
One of the main hypotheses of the WWT, is the assumption of random phases and amplitudes. When the nonlinearity is strong, phases might be strongly correlated and spectrum could be guessed by invoking the critical balance (CB) assumption where linear and nonlinear-transfer terms in (1) are balanced. It follows that the spectrum is independent of the flux ($y=0$). Such an assumption, first evoked in the context of water waves, led to the so-called Phillips spectrum Phillips (1958). Here, unlike water waves where the CB assumption fixes all exponents, we still need to fix the scaling exponent $x$. For GPE, the CB spectrum was first introduced in Proment et al. (2009) using the GPE (1) written in Fourier space
$i\frac{\partial\hat{\psi}_{\bf k}}{\partial t}=\omega_{k}\hat{\psi}_{\bf k}+\beta\sum_{123}\hat{\psi}_{1}^{*}\hat{\psi}_{2}\hat{\psi}_{3}\delta^{{\bm{k}}1}_{23},\,{\rm with}\,\omega_{k}=\alpha k^{2},\,k=|{\bf{k}}|,$ (6)
where $\sum_{123}$ stands for a sum over all wave vectors ${\bf{k}}_{1},{\bf{k}}_{2},{\bf{k}}_{3}$, $\psi_{\bf i}=\psi({\bf{k}}_{i})$, and $\delta^{{\bm{k}}1}_{23}$ is a Kronecker-$\delta$: 11 if ${\bf{k}}+{\bf{k}}_{1}={\bf{k}}_{2}+{\bf{k}}_{3}$, 0 otherwise. Note that for both WWT and CB, the leading-order nonlinear effect in (6) is provided by the diagonal terms in the sum, i.e., the ones with ${\bf{k}}={\bf{k}}_{2},{\bf{k}}_{1}={\bf{k}}_{3}$ and ${\bf{k}}={\bf{k}}_{3},{\bf{k}}_{1}={\bf{k}}_{2}$. They lead to a nonlinear frequency shift $\omega_{k}\to\omega_{k}+2\beta N$, but do not affect the evolution of $|\hat{\psi}_{{\bm{k}}}|$ and, therefore, do not lead to any spectral redistribution. They can be eliminated by passing into the rotating frame, $\hat{\psi}_{\bf k}\to e^{-2i\beta Nt}\hat{\psi}_{\bf k}$. Applying the CB argument to (6), we find that $\omega_{k}|\hat{\psi}_{\bf k}|\sim\beta|\hat{\psi}_{\bf k}|^{3}(k/k_{\rm cor})^{6}$ where $k_{\rm cor}$ is the correlation distance in the $k$-space (it is vanishingly small in the WWT). The $(k/k_{\rm cor})^{6}$-term estimates the number of modes interacting with mode $k$. With $n_{\bf k}\sim|\hat{\psi}_{\bf k}|^{2}/k_{\rm cor}^{3}$, this leads to $n_{\bf k}\sim\frac{\alpha}{\beta k}\frac{k_{\rm cor}^{3}}{k^{3}}.$ In the limit $L\to\infty$, this expression should be independent of $L$, which arises if we postulate $k_{\rm cor}\sim\alpha/\beta$. This finally gives
$n_{\bf k}\sim\left(\frac{\alpha}{\beta k}\right)^{4}$ (7)
i.e. Eq.(4) with $x=4$.
Numerical simulations.— We perform numerical simulations of the forced-dissipated GPE using the standard massively-parallel pseudo-spectral code FROST Krstulovic (2020) with a fourth-order Exponential Runge-Kutta temporal scheme (see Zhu et al. (2023)). We set $L=2\pi$, $\alpha=\beta=1$, and use grids of $N_{p}^{3}$ collocation points, with $N_{p}=512$. We add a forcing term $F_{\bf k}(t)$ and a dissipation term $D_{\bf k}\widehat{\psi}_{\bf{k}}(t)$ to the RHS of the Fourier-space GPE (6). To obtain inverse cascade the forcing term is supported on a narrow band at high $k$’s around the forcing wavenumber $k_{\rm f}$ for $k\in[k_{\rm f}-1,k_{\rm f}+1]$, and it is given by $\mathrm{d}F_{\bf k}(t)=f_{0}\mathrm{d}{W}_{\bf k}$, where $W_{\bf k}$ is the Wiener process, and the parameter $f_{0}$ controls the amplitude of the forcing. Dissipation is of the form $D_{\bf k}=D_{\rm L}k^{-r}$ at the largest scales ($0<k\leq k_{\rm L}\ll k_{\rm f}$), and $D_{\bf k}=D_{\rm R}k^{s}$ at the very small scales less than the forcing scales ($k_{\rm f}+1<k\leq k_{\rm max}$). Additionally, the condensate mode ($k=0$) is dissipated at a constant rate $D_{\bf 0}$. We optimize the parameters of forcing and dissipation to enlarge the inertial range for a fixed resolution, ensuring that the simulations that are well-resolved and minimizing bottlenecks at the dissipation scales. All the numerical parameters are listed in the Supplemental Material (SM). We perform simulations with different values of the forcing strength $f_{0}$ in order to study the transition from the weak to the strong WT regimes. The $k$-space particle flux $Q(k)$ is computed directly from the RHS of the GPE (6) as in Zhu et al. (2023). We evaluate $Q_{0}$ (the rate of particle loss per unit volume and unit time through low-k dissipation) by measuring the value of $Q(k)$ in the inertial range where it is approximately constant. Note that $f_{0}$ and $|Q_{0}|$ are monotonically related, so we discuss our results in terms of the strength of the particle flux. Finally, our numerical measurements are performed when a statistically steady state is reached.
Figure 1(a) shows the stationary waveaction spectrum $n(k)$ for values of the flux $Q_{0}$ varying over almost five order of magnitude. The respective spectra compensated by the KZ prediction (5) and the wavenumber compensated by the healing length $\xi=1/\sqrt{N}$ are shown in Fig.1(b).
Figure 1: (a) Stationary spectra for different strengths. of the particle flux $|Q_{0}|$ (b) Same spectra compensated by the KZ prediction.
We observe that for lower values of $|Q_{0}|$ there is almost perfect agreement with the WWT KZ spectrum prediction (5) – with both the slope and the prefactor constant of this solution confirmed (as previously seen in Zhu et al. (2023)). Naturally, the range at which this agreement holds shrinks when the flux $|Q_{0}|$ increases, and the weak nonlinearity assumption fails at $k\xi\lesssim 1$. This is natural because the healing length $\xi$ is defined as the scale where the linear and the nonlinear terms in the GPE are of a similar size. As scales $k\xi\lesssim 1$, the $k^{-4}$-CB scaling (7) appears when the flux is moderately strong such that $k_{\xi}=1/\xi$ still lays within the inertial range. Note, however, that CB balance can only be observed as an intermediate state for moderately high fluxes, since equating the KZ and CB spectra yields a crossover wave number increasing with the flux as $|Q_{0}|^{1/5}$, which is a paradoxical prediction of an increase of the WWT range of scales with an increased flux. Therefore, CB is not the ultimate strong wave turbulence regime. Indeed, for the highest particle flux, we observe a very steep spectrum at low $k$’s and close to a $k^{-2}$-spectrum at large wavenumbers. We will later relate this behavior to the emergence of a quasi-condensate with a thermal component induced by the forcing.
In order to quantitatively study the transition from the WWT to the strong wave turbulence, we make use of the spatio-temporal Fourier transform (STFT) of the wavefunction, which is obtained by performing the Fourier transform in both physical-space and in time over a finite (but sufficiently large) time window. STFT has become a standard tool in wave turbulent systems for checking if the system is in the weakly nonlinear regime: in this case STFT is narrowly concentrated in the proximity of the linear-wave dispersion relation $\omega_{k}=\alpha k^{2}$, i.e. the width of the frequency broadening of STFT spectrum at a fixed $k$, $\delta\omega(k)$, must remain much less than $\omega_{k}$. Moreover, the CB and strong wave turbulence regimes are characterized by $\omega_{k}\sim\delta\omega_{k}$ and $\omega_{k}<\delta\omega_{k}$, respectively.
Representative examples of the STFT spectral plots are shown in Figs .2(a)–(c) and the measurements of the nonlinear frequency broadening $\delta\omega(k)$ (defined here as the width corresponding to $80\%$ of the area under the curve) are in Fig.1(d).
Figure 2: (a)–(c): Representative STFT spectra for the weak, intermediate and strong fluxes; the dashed-dotted vertical lines are at $k=1/\xi$. (d): nonlinear frequency broadening as a function of $k$ for the cases shown in (a)–(c), (the straight line is $\delta\omega=\omega_{k}$. (e) Strong flux case: spectrum decomposition into the condensate and thermal components.
We can see that in both the weak (a) and the intermediate flux cases (b), the spectrum is distributed closely to the dispersion relation $\omega=\alpha k^{2}+2N$, with $2N$ being the nonlinear shift. The nonlinear broadening displayed in Fig.2.d shows that for the lowest flux case the weak nonlinearity condition for applicability of weak WT, $\delta\omega(k)\ll\omega_{k}$ is realized over a broad inertial range of wavenumebers between the large-scale dissipation and the forcing scales, $3\lesssim k\lesssim 100$, the range where the inverse cascade KZ prediction is clearly observed. For the intermediate flux case, the range where $\delta\omega(k)\ll\omega_{k}$ is also present, but now in a narrower range, $30\lesssim k\lesssim 100$. KZ spectrum is still seen in this range. On the left of it, in range $10\lesssim k\lesssim 30$, we observe $\delta\omega(k)\sim\omega_{k}$, i.e. the CB state. Respectively, the CB spectrum is indeed realised in this range. The transition between the KZ and the CB scaling occurs approximately at the healing length wavenumber, $k_{\xi}\sim 30$. The strong WT case deserves more explanation.
Emergence of condensate and acoustic turbulence.— When the flux is very intense, the CB assumption is not longer fulfilled and the system transits to a new regime of strong wave turbulence, exhibiting a very steep spectrum at low $k$’s (Fig. 1). For this case, the STFT in Fig. 3.c displays a behavior completely different from a broadened $\omega_{k}=\alpha k^{2}+2N$ curve. This state is characterized by the emergence of a strong coherent condensate component around $k=0$. The condensate affects, nonlocally in the $k$-space, the dynamics of waves with higher wavenumbers, thereby invalidating the locality property used by both the WWT and the CB theories. The effect of the condensate takes the form of making the wave perturbations behave as sound characterised by the well-known Bogoliubov dispersion relation,
$\omega_{B}(k)=\beta|\psi_{0}|^{2}\pm\sqrt{2\alpha\beta|\psi_{0}|^{2}k^{2}+\alpha^{2}k^{4}},$ (8)
where $|\psi_{0}|^{2}=|\hat{\psi}_{{\bf k=0}}|^{2}$ is the condensate density (which is $\approx N$ for a strong condensate). The Bogoliubov dispersion relation is clearly observed in the STFT in Fig. 3.c, including the nonlinear frequency gap.
Like in the CB regime, the Fjørtoft dual cascade argument does not work in this regime. In the acoustic regime, there is no local inverse cascade, and most of $N$-flux is now toward the high $k$’s, whereas the low-$k$ range distribution is necessarily close to thermodynamic equilibrium (i.e., the role of the IR dissipation is negligible). The classical picture for such equilibrium is a condensate distribution $N_{0}\delta{({\bf{k}})}$ superimposed with thermal fluctuations in the form of the energy equipartition of the modes, the so-called Bogoliubov spectrum $n_{B}(k)\propto T/k^{2}$ with constant $T$, which has the meaning of temperature. In a out-of-equilibrium three-wave regime (such as in a presence of a condensate), the flux $Q$ is proportional to the second power of $n_{k}$, and thus $T\sim|Q_{0}|^{1/2}$, which gives
$n_{B}(k)\propto\frac{|Q_{0}|^{1/2}}{k^{2}}.$ (9)
Note that the Bogoliubov $k$-scaling (9) coincides with the Rayleigh-Jeans energy equipartition scaling for the weakly interacting Bose gas governed by the four-wave mixing process. Finally, in a finite box, the condensate distribution takes the form of a ”widened” $\delta$-function, i.e., a distribution sharply peaked around $k=0$ and rapidly decaying for $k\neq 0$.
The presence of such a condensate corresponds to the short horizontal line at $\omega\approx 1.34N$ in the STFT of Fig. 2.c, and enhanced in the inset of pannel (e). Note that for pure uniform condensate, the STFT spectrum would be concentrated at a point $(k,\omega)=(0,N)$. In our case the condensate spectrum is spread in $k$, i.e. the condensate is non-uniform in the physical space. Also, its frequency is shifted upward with respect to the uniform condensate case. Both of these features can be linked to the finite size of the computational box. One can take advantage of the fact that the wave and the condensate components in the large flux case are clearly separated on the STFT plot, and calculate the spectra of the waves and the condensate separately. To do that we band-pass filter the frequency spectrum in a narrow range $1.30N<\omega<1.38N$ and call it condensate, while the rest of the spectrum we call waves. Then, performing the inverse Fourier transform in $\omega$ we get the $k$-spectra shown in Fig. 2(d), showing a thermodynamic Bogoliubov spectrum of waves over an extended $k$-range and a very peaked low-$k$ condensate’s spectrum decaying approximately as $k^{-7}$.
Let us clarify about the role of vortices in our system for different forcing intensities. Vortices are associated with incompressible kinetic energy, which arises from the Helmholtz decomposition of the regularized velocity field into solenoidal and potential parts. Fig. 3(a) displays the energy components for various values of $|Q_{0}|$: incompressible and compressible kinetic energies, internal energy, and quantum energy. For low $|Q_{0}|$, quantum energy is dominant and internal energy is minimal, consistent with weak nonlinearity. Compressible and incompressible energies are equal, and each is half the quantum energy. The presence of incompressible kinetic energy indicates vortices, but these are ”ghost” vortices without dynamical significance—they appear even in purely linear fields Berry and Dennis (2000). Interestingly, increasing $|Q_{0}|$ reduces incompressible kinetic energy to about 6–8% in the strongest forcing cases, reflecting a sparse distribution of small vortex loops (with radii of the order of the healing length). This state contrasts with strong superfluid hydrodynamic turbulence states consisting of polarized tangles of fully formed quantized vortex lines of all sizes Polanco et al. ; Fischer and Bradley (2025). In our system, strong vortex lines shrink and annihilate due to effective friction from the intense acoustic component. (Persistent vortex tangles would require sound damping.) Thus, the strong turbulence state reported here is dominated by interacting acoustic waves rather than dynamically significant hydrodynamic vortices.
We will now discuss the out-of-equilibirum equation of state Dogra et al. (2023); Zhu et al. (2024), which relates the global amplitude of the waveaction spectrum to the input flux, in our case $|Q_{0}|$. In practice, we fit our data with $n_{k}=Ak^{-x}$ in the different observed ranges. The exponents $x$ was found to be very close those discussed earlier, and the amplitude $A$ is strongly dependent on $Q_{0}$. Readers can find the fitting method and fitting exponents $x$ in the SM. Figure 3(b) shows the plot of the spectrum’s amplitude $A$ vs. the particle flux in the inertial range $|Q_{0}|$ for our runs covering the whole range of particle fluxes, from low to high (the flux functions $Q(k)$ for all the runs are shown in the SM). For low-flux runs, we observe $A\sim|Q_{0}|^{1/3}$ in agreement with the weak four-wave WT. For intermediate fluxes, the spectrum at the high-$k$ part remains weak and still follows $A\sim|Q_{0}|^{1/3}$. The low-$k$ part, which exhibits the CB exponent (see SM), displays some dependence on the flux that could be fitted by $A\sim|Q_{0}|^{2/3}$. Such an exponent evokes some type of hydrodynamic Kolmogorov turbulence, but we lack a theoretical explanation and scaling range is too narrow. Finally, for high-flux runs, we only measure $A$ for the large-$k$ thermodynamic part because the low-$k$ condensate spectrum range is too short for a reliable fitting. Here we see $A=T\sim|Q_{0}|^{1/2}$, which fulfills our prediction for the temperature scaling of the three-wave process.
Figure 3: (a): Percentages of energy components in terms of flux for the incompressible kinetic energy $E_{\rm kinI}$, compressible kinetic energy $E_{\rm kinC}$, quantum energy $E_{\rm q}$, and internal energy $E_{\rm int}$, respectively; (b): “Equation of state” plot covering the weak, intermediate, and strong flux regimes.
Conclusions.— In this Letter, we studied the stationary inverse cascade states in BEC turbulence for a large range of forcing strengths (or equivalently particle fluxes). We showed that the predictions of the WWT theory are very robust; namely, the KZ spectrum scaling is observed for fluxes spanning over more than four orders of magnitude. At stronger fluxes, the waveaction spectrum and the STFT spectrum indicate the presence of the Critical Balance state in which the nonlinear frequency broadening $\delta\omega(k)$ is close to the linear wave frequency $\omega_{k}$ in a range of wavenumbers $k$. In the simulations with the strongest flux, we observe states close to low-temperature thermodynamic equilibria: with a prominent condensate component sharply peaked near $k=0$ and over-condensate fluctuations with spectra close to the thermal Bogoluibov spectrum $n_{k}\sim 1/k^{2}$. Our complete set of simulations has allowed to determine the a rich equation of state for the 3D inverse cascade BEC turbulence.
Acknowledgements.
This work was funded by the Simons Foundation Collaboration grant Wave Turbulence (Award ID 651471). This work was provided with computing HPC and storage resources of IDRIS and CINES under the allocations A0102A12494 and A0172A14637 made by GENCI on the Jean Zay (SKL partition) and Adastra (GENOA partition) supercomputers. This work was granted access to the OPAL infrastructure from Université Côte d’Azur, supported by the French government, through the UCAJEDI Investments in the Future project managed by the National Research Agency (ANR) under Reference No. ANR-15-IDEX-01.
Appendix A Numerical parameters
case $f_{0}$ $D_{\rm L}$ $r$ $k_{\rm L}$ $D_{\rm R}$ $s$ $D_{\bf 0}$
0 $10^{-4}$ 1 0.5 6 $4.29\times 10^{-26}$ 6 $10^{3}$
1 $2\times 10^{-4}$ 2 0.5 6 $2.54\times 10^{-30}$ 7 $10^{3}$
2 $4\times 10^{-4}$ 2 0.25 7 $1.50\times 10^{-34}$ 8 $10^{3}$
3 $8\times 10^{-4}$ 4 0.25 8 $8.89\times 10^{-39}$ 9 $10^{3}$
4 $1.6\times 10^{-3}$ 8 0.25 8 $5.26\times 10^{-43}$ 10 $10^{3}$
5 $3.2\times 10^{-3}$ 10 0.25 8 $3.11\times 10^{-47}$ 11 $10^{3}$
6 $1.2\times 10^{-2}$ 16 0.25 8 $1.63\times 10^{-55}$ 13 $10^{3}$
7 $2.5\times 10^{-2}$ 70 0.5 10 $1.63\times 10^{-55}$ 13 $10^{3}$
8 $5\times 10^{-2}$ 250 0.75 11 $9.75\times 10^{-63}$ 15 $10^{4}$
9 0.2 16 0.125 8 $9.75\times 10^{-63}$ 15 $10^{4}$
10 0.5 20 0.125 8 $9.75\times 10^{-63}$ 15 $10^{4}$
Table S1: Parameters for GPE simulations.
All essential numerical parameters for the numerical simulation runs are given in Table S1. To systematically study the transition from weak wave turbulence to strong turbulence, we fix the size of the computing domain in Fourier space with $k_{\rm max}=173$, and the forcing position with $k_{\rm f}=125$ for all the runs.
Appendix B Flux spectra
We plot the flux spectra $|Q(k)|$ for all the runs in Figure S1 for $k\leq k_{\rm f}$.
Figure S1: Spectra flux $|Q(k)|$. The particle flux $|Q_{0}|$ is measured in the inertial ranges with quasi-constant values.
Appendix C Method to generate the EoS figure
Suppose the spectrum is of the form $n=A\,k^{-x}$, we compute the values of $A$ and $x$ by fitting the power-law range of the numerically observed spectrum with $\ln{n}=\ln{A}-x\ln{k}$. Figure S2 presents how we fit spectra for the strong turbulence cases.
Figure S2: Fitting spectra of the thermal components for case 9 and case 10: (a) Spectra compensated by the expected power-law solution (Bologliubov spectrum $n_{\rm B}\sim k^{-2}$); (b) Original spectra with the power-law fits superposed.
The fitting exponents corresponding to Figure 3 of the main text is shown in Figure S3.
Figure S3: The fitting exponents correspond to Figure 3 of the main text. The theoretical predictions for weak wave turbulence, critical balance, and the Bogoliubov spectrum are 7/3-7/3, 4-4, and 2-2, respectively.